# Rad4XCNN: a new agnostic method for post-hoc global explanation of CNN-derived features by means of radiomics

### v1

@@ Abstract
In the last years, artificial intelligence (AI) in clinical decision support systems (CDSS) played a key role in harnessing machine learning and deep learning architectures. Despite their promising capabilities, the lack of transparency and explainability of AI models poses significant challenges, particularly in medical contexts where reliability is a mandatory aspect. Achieving transparency without compromising predictive accuracy remains a key challenge. This paper presents a novel method, namely Rad4XCNN, to enhance the predictive power of CNN-derived features with the interpretability inherent in radiomic features. Rad4XCNN diverges from conventional methods based on saliency map, by associating intelligible meaning to CNN-derived features by means of Radiomics, offering new perspectives on explanation methods beyond visualization maps. Using a breast cancer classification task as a case study, we evaluated Rad4XCNN on ultrasound imaging datasets, including an online dataset and two in-house datasets for internal and external validation. Some key results are: i) CNN-derived features guarantee more robust accuracy when compared against ViT-derived and radiomic features; ii) conventional visualization map methods for explanation present several pitfalls; iii) Rad4XCNN does not sacrifice model accuracy for their explainability; iv) Rad4XCNN provides global explanation insights enabling the physician to analyze the model outputs and findings. In addition, we highlight the importance of integrating interpretability into AI models for enhanced trust and adoption in clinical practice, emphasizing how our method can mitigate some concerns related to explainable AI methods.
@@ 1 Introduction
The use of computer-aided tools based on Artificial Intelligence (AI) methods has increased significantly. These tools utilize machine learning and deep learning architectures in many types of Decision Support Systems (DSS). A particular class of DSS is represented by Clinical decision support systems (CDSS), implemented to support activities in critical healthcare processes. Although data-driven methods play a crucial role in CDSSs development, their use in medicine still harbors many pitfalls. New deep learning methods and the growing data availability have enabled the development of powerful but uninterpretable CDSS, and the field of Explainable AI (XAI) has gained considerable attention. Regulatory authorities have addressed the problem of explainability Kundu (2021); Arrieta et al. (2020). The US Federal Trade Commission emphasizes that AI tools should be transparent, explainable, fair, and empirically based while promoting accountability Smith and Director (2020). The European Parliament has adopted the General Data Protection Regulation (GDPR), in which meaningful explanations of the logic involved are declared mandatory when automated decision-making is performed Guidotti et al. (2018). The lack of transparency makes both doctors and patients skeptical about these new technologies. Opaque AI systems can impair the doctor-patient relationship and jeopardize patient trust Amann et al. (2020). For this reason, scientific research is working on making AI transparent and/or explainable Combi et al. (2022). In medical contexts, several aspects can affect the model reliability and demand explainability Holzinger (2016); Holzinger et al. (2019). Consequently, methods that enable transparency and explainability can help validate the models, improve the knowledge domain, and increase the real-world usability of the systems.
For these reasons, although it is not yet fully accepted Bornstein (2016); Ghassemi et al. (2021); McCoy et al. (2022); London (2019), explainability is increasingly becoming a requirement that these systems should fulfill Jovanović and Schmitz (2022).
@@
Deep learning architectures have demonstrated remarkable capabilities in extracting intricate patterns and features from huge datasets, enabling unprecedented predictive performance in various fields, including medical image analysis. Although deep features achieve impressive accuracy, they ignore the question of explainability and several approaches were proposed for their post-hoc explanation. These approaches focus mainly on the concept of visualization maps that have shown several limitations regarding their reliability Cerekci et al. (2024); Prinzi et al. (2024a); Zhang et al. (2021).
Radiomics Gillies et al. (2016); Lambin et al. (2017) is a new alternative approach to introduce explainability within the feature extraction process in medical imaging Prinzi et al. (2024b). In Radiomic, features are calculated by appropriate mathematical formulas applied to the greyscale histogram, to texture-defining matrices or ROIs shape quantifiers. As a result, the meaning of each radiomic feature is well known and meaningful clinical conclusions can be drawn through model explanation Prinzi et al. (2024c).
Although radiomic features have the great property of inherent explainability, their predictive power seems considerably weaker than deep features, as has been shown in several works Lisson et al. (2022); Sun et al. (2020); Truhn et al. (2019); Wei (2021). Again, the dilemma concerning the trade-off between explainability and accuracy arises. Hence, the use of deep features would ensure highly performing but poorly interpretable models, in contrast, radiomic features would ensure less accurate but explainable models.
@@
In this study, we attempt to combine the predictive power of deep features with the intrinsic explainability characteristic of radiomic features. To this aim, state-of-the-art architectures such as ResNet, DenseNet, and Vision Transformer (ViT) for deep feature extraction were employed. Simultaneously, a radiomic workflow to extract interpretable signatures was implemented. We performed a comparative analysis between these two feature sources in terms of predictive performance and explainability. Moreover, the new method Rad4XCNN proposed in this work combines the predictive power of the deep features while exploiting the inherent interpretability of the radiomic features. In particular, Rad4XCNN doesn’t introduce constraints or bottlenecks that can harm accuracies in deep architecture training while explaining CNN-derived features by means of radiomics. In addition, it provides new perspectives on the development of explanation methods not solely relying on visualization maps Papanastasiou et al. (2023). A breast cancer classification task was proposed as a case study to evaluate the effectiveness of Rad4XCNN: this application reflects the challenges encountered in real-world clinical scenarios. To this end, we acquired two in-house datasets, one for training and internal testing and the other for external validation.
@@
  • the use of deep architectures for breast cancer classification in ultrasound (US) imaging exploiting an online dataset, and two in-house datasets for internal and external tests;
@@ 2 Background
Explainable AI
@@
Numerous definitions of explainable AI are not fully standardized yet Longo et al. (2024). For this article, it is essential to delineate certain concepts. A model is intrinsically interpretable when its transparent structure allows an understanding of its decision process. For this reason, a transparent algorithm doesn’t require explanation methods for its interpretation. From an algorithmic point of view, Logistic Regression, Decision Tree, and Naive Bayes are transparent. Neural Networks (NNs), Support Vector Machines (SVMs), and Tree Ensemble (TE) are defined as black-boxes and require explanation methods for their introspection. An explanation can be global and local. A global explanation aims to explain a model’s behavior across its entire dataset. A local explanation aims to explain the prediction of a specific dataset sample (i.e. patient). To an intelligible feature it is possible to associate a human-understandable meaning. Radiomic features are considered intelligible. Conversely, deep features (e.g. extracted via neural network architectures) are not intelligible. We can call features extracted via convolutional neural networks CNN-derived, and extracted via Vistion Transformer as ViT-derived.
@@
The combination of intelligible features and transparent algorithms makes the whole system intrinsically interpretable for two reasons: 1) it is possible to calculate the impact of the feature for model decision; 2) it is possible to correlate and validate the model’s findings comparing it with the clinical literature. When black-box algorithms are used and the features are intelligible, explanation methods allow for global and local explanations. Typically, these methods are called post-hoc algorithms because they are applied after model training. Model explainability becomes much more complicated when not-intelligible features are used: model findings cannot be correlated with clinical literature because the meaning of each feature is unknown (as happens with learned features in deep architectures). A method is defined agnostic when can be applied to any algorithm and architecture. For classification tasks, an XAI method can be class-dependent when explanations are given for each class separately, conversely, they are defined as class-independent.
Explainability in Deep Architectures
@@
Convolutional-based neural networks have emerged as a standard for deep feature extraction in medical images. The explainability of these architectures is mainly focused on the concept of saliency maps, implemented to highlight the image areas most informative for the prediction Itti et al. (1998); Mamalakis et al. (2023). In Simonyan et al. (2013) the gradient computation of the class score concerning the input image is exploited to visualize the activation map specific to a given image and a particular class. Integrated gradients are proposed in Sundararajan et al. (2017) for a more robust explanation through activation maps, adding the constraint of respecting the ’sensitivity’ and the ’implementation invariance’ principles/axioms. In Zeiler et al. (2011) and Zeiler and Fergus (2014) a network using deconvolution is proposed to visualize the convolutional network and in Springenberg et al. (2014) a variant of the deconvolution approach with guided back-propagation is proposed to visualize the features learned from CNN. The most popular algorithms exploiting saliency belong to the category of ’class activation maps’ (CAM) Zhou et al. (2016). GradCAM Selvaraju et al. (2017) has been introduced after CAM, to overcome the limitation of being able to visualize only the last layer and only particular CNN architectures. Despite many other methods being proposed later, GradCAM remains one of the most widely used. Lately, methods called ’gradient-free’ are also gaining wide popularity to overcome some limitations for computing gradients in very-deep architectures, and overcome the problem of anomalous behavior in incorrect prediction scenarios Muhammad and Yeasin (2020).
@@
Saliency maps are mainly used to provide local explanations and it is not clear how these tools can be used for global explanation. Furthermore, it is important to emphasize that criticisms have been raised about their use Prinzi et al. (2023a). For example, poorly localized, and spatially blurred visualization was found in some cases Oh et al. (2020); Signoroni et al. (2021). Moreover, it has been shown that different saliency map computation techniques can produce conflicting results Prinzi et al. (2024a); Zhang et al. (2021).
Radiomic Workflow
@@
Radiomics is gaining prominence for extracting features from radiological images. It operates on the premise that some pixel-wise patterns are hidden from the human eye. For this reason, it is defined as a workflow aimed to convert standard-of-care medical images into high-dimensional data Gillies et al. (2016). In contrast to deep features, radiomic features are computed through mathematical formulas applied to the images and quantify some characteristics such as texture/pattern and statistical measurement. For this reason, radiomic features are also called hand-crafted, counterposed to the learned extracted by neural networks. Radiomic flow has several advantages over deep extraction. While deep extraction requires large datasets for model training, radiomic feature extraction can be performed with small datasets Wei (2021). Moreover, shallow learning algorithms prove to be well-suited for classifying radiomic data, widely recognized as more appropriate for training with small datasets. The main advantage lies in the inherent interpretability of radiomic features Militello et al. (2023); Prinzi et al. (2024b): it is well known the meaning each feature expresses. The use of shallow learning algorithms and intrinsically interpretable features enables the local and global introspection of the predictive models.
Our Rationale
Despite in some applications the use of transparent models can be sufficient Rudin (2019), the design of a transparent model can introduce a level of complexity Chen et al. (2022): interpretable features are typically less informative than deep features. In addition, the use of transparent algorithms may fail when the relationships between the data are complex and nonlinear. As a consequence, the use of transparent-by-design models can lead to poor performance. For this reason, the eternal conflict between the explainability and accuracy trade-off comes into play.
The literature indicates the need for new methods that deviate from those based on visualization maps Papanastasiou et al. (2023). Some attempts have been made by Paul et al. (2019); Wang et al. (2019); Chen et al. (2016) although there is no explicit reference to XAI. Paul et al. (2019) evaluated how radiomic model performance varies by replacing a correlated deep feature in the radiomic signature. However, regarding the performed interpretable feature extraction, the Image Biomarkers Standardization Initiative (IBSI) guidelines Zwanenburg et al. (2020) had not yet been proposed at that time. As a consequence, assessing the reproducibility of the study and the resulting impact of the study can be challenging. In addition, the authors do not use state-of-the-art deep extraction architectures such as ResNet, DenseNet, and ViT, whose depth is quite different from the VGG employed. Furthermore, they agree with an important limitation regarding the fact that a single slice was used to extract deep features in a volumetric image, whereas semantic information was generated from multiple slices. Above all, in contrast to our idea, they modify the input signature by replacing semantic features with deep ones, introducing a kind of constraint for model training. Wang et al. (2019) attempts to exploit the radiologist’s knowledge to qualitatively and visually correlate CNN-based feature maps with semantic features related to the investigated domain. However, these radiological features are operator-dependent, as is their correlation with the feature maps. As a result, their proof-of-concept is not generalizable to any context or domain. A similar idea was proposed in Chen et al. (2016), to effectively bridge the gap between deep features and multiple clinical semantic features. The model aims to yield semantic rating scores to support a deeper analysis for either clinical diagnosis or educational purposes. More recently, Liu et al. (2023) tried to train deep models to grade informative semantic features related to the analyzed task rather than directly predicting the disease malignancy. Both papers lack formalism regarding the terminology related to explainable AI (i.e., the characteristics of the method in terms of global-local explanation, applicability, class dependencies, etc.) In general, in the attempt to make the predictive process transparent, the cited articles, employ semantic features subject to inter-operator variability or modify the training process, which could harm the established predictive performance of deep models.
In this paper, we present Rad4XCNN, a method that integrates the predictive strengths of deep features with the interpretability offered by radiomic features. The method does not introduce any constraints in the training that inhibit its accuracy. Considering the development of new explanation methods not solely based on visualization maps as a major challenge Papanastasiou et al. (2023), Rad4XCNN offers an innovative approach to explain deep architectures and provides a global explanation of deep networks, an aspect often overlooked in current research.

### v2

@@ Abstract
-Background and Objective: In recent years, machine learning-based clinical decision support systems (CDSS) have played a key role in the analysis of several medical conditions. Despite their promising capabilities, the lack of transparency in AI models poses significant challenges, particularly in medical contexts where reliability is a mandatory aspect. However, it appears that explainability is inversely proportional to accuracy. For this reason, achieving transparency without compromising predictive accuracy remains a key challenge. Methods: This paper presents a novel method, namely Rad4XCNN, to enhance the predictive power of CNN-derived features with the inherent interpretability of radiomic features. Rad4XCNN diverges from conventional methods based on saliency maps, by associating intelligible meaning to CNN-derived features by means of Radiomics, offering new perspectives on explanation methods beyond visualization maps. Results: Using a breast cancer classification task as a case study, we evaluated Rad4XCNN on ultrasound imaging datasets, including an online dataset and two in-house datasets for internal and external validation. Some key results are: i) CNN-derived features guarantee more robust accuracy when compared against ViT-derived and radiomic features; ii) conventional visualization map methods for explanation present several pitfalls; iii) Rad4XCNN does not sacrifice model accuracy for their explainability; iv) Rad4XCNN provides a global explanation enabling the physician to extract global insights and findings. Conclusions: Our method can mitigate some concerns related to the explainability-accuracy trade-off. This study highlighted the importance of proposing new methods for model explanation without affecting their accuracy.
@@ 1 Introduction
The use of computer-aided tools based on Artificial Intelligence (AI) methods has increased significantly. These tools utilize machine learning and deep learning architectures in many types of Decision Support Systems (DSS). A particular class of DSS is represented by Clinical decision support systems (CDSS), implemented to support activities in critical healthcare processes. Although data-driven methods play a crucial role in CDSSs development, their use in medicine still harbors many pitfalls. New deep learning methods and the growing data availability have enabled the development of powerful but uninterpretable CDSS, and, for this reason, the field of Explainable AI (XAI) has gained considerable attention. Regulatory authorities have discussed the lack of explainability Kundu (2021); Arrieta et al. (2020). The US Federal Trade Commission emphasizes that AI tools should be transparent, explainable and fair Smith (2020). The European Parliament has adopted the General Data Protection Regulation (GDPR), in which meaningful explanations of the logic involved are declared mandatory when automated decision-making is performed Guidotti et al. (2018). Finally, with the Regulation (EU) 2024/1689 of the European Parliament, on July 2024 was approved the AI Act, establishing harmonized rules on the use of artificial intelligence-based systems European Community (2024). In this perspective, the AI Act aims to achieve its stated policy objectives with a focus on transparency and human oversight Panigutti et al. (2023).
From a practical point of view, the lack of transparency makes both doctors and patients skeptical about these new technologies. Opaque AI systems can impair the doctor-patient relationship and jeopardize patient trust Amann et al. (2020). For this reason, scientific research is working on making AI transparent and/or explainable Combi et al. (2022). In medical contexts, several aspects can affect the model reliability and demand explainability Holzinger (2016); Holzinger et al. (2019). Consequently, methods that enable transparency and explainability can help validate the models, improve the knowledge domain, and increase the use of these systems in the real world. For these reasons, although it is not yet fully accepted Bornstein (2016); Ghassemi et al. (2021); McCoy et al. (2022); London (2019), explainability is increasingly becoming a requirement that these systems should fulfill Jovanović and Schmitz (2022).
@@
Deep learning architectures have demonstrated remarkable capabilities in extracting intricate patterns and features from huge datasets, enabling unprecedented predictive performance in various fields, including medical image analysis. Although deep features achieve impressive accuracy, the question of explainability is ignored and several approaches were proposed for their post-hoc explanation. These approaches focus mainly on visualization maps, which have shown several limitations regarding their reliability. As an example, applying different saliency map methods can result in different explanations Cerekci et al. (2024); Prinzi et al. (2024a); Zhang et al. (2022). Furthermore, while saliency maps provide a form of local explanation, it has been shown that a global explanation is necessary for clinical model validation Prinzi et al. (2023a).
Radiomics Gillies et al. (2016); Lambin et al. (2017) is a new alternative approach to introduce explainability within the feature extraction process in radiological imaging Prinzi et al. (2024b). Radiomic uses mathematical formulas to analyze grayscale histograms, ROI shapes, or texture-defining matrices. Consequently, each radiomic feature’s significance is well known, and meaningful clinical conclusions can be drawn through model explanation Prinzi et al. (2024c). In addition, radiomics is more suitable for training in small dataset scenarios. In fact, model training relies on algorithms designed for tabular data analysis, which are more appropriate for training with limited data An et al. (2021); Traverso et al. (2018). Although radiomic features have the great properties of training in small data scenarios and inherent explainability, their predictive power seems considerably weaker than deep features, as has been shown in several works Lisson et al. (2022); Sun et al. (2020); Truhn et al. (2019); Wei (2021). As a result, deep learning architectures are implemented when accuracy is the metric to maximize Truhn et al. (2019).
Considering that the feature extraction process from medical images represents the main step to implement high accuracy and explainability, several studies analyzed the interpretable capability of radiomic features Varriano et al. (2022) and the differences between learned features and radiomic features Rundo and Militello (2024). The dilemma concerning the trade-off between explainability and accuracy arises van der Veer et al. (2021a). Hence, the use of deep features would ensure highly performing but poorly explainable models, in contrast, radiomic features would ensure less accurate but interpretable models.
@@
In this study, we combine the predictive power of deep features with the intrinsic explainability characteristic of radiomic features. To this aim, state-of-the-art architectures such as ResNet, DenseNet, and Vision Transformer (ViT) for deep feature extraction were employed Rahman et al. (2024). Simultaneously, a radiomic workflow to extract interpretable signatures was implemented. We performed a comparative analysis between these two feature sources in terms of predictive performance and explainability. Moreover, we propose Rad4XCNN, a new method leverages deep learning for enhanced performance while exploiting the inherent interpretability of radiomic features. In particular, from a methodological perspective, Rad4XCNN doesn’t introduce constraints or bottlenecks that can harm accuracies in deep architecture training Alvarez Melis and Jaakkola (2018); Elbaghdadi et al. (2020) while it explains CNN-derived features by means of radiomics. From a clinical perspective, Rad4XCNN uses quantitative radiomic features for explanations, introduces a form of global explanation, enriches image interpretation with intelligible features, and contributes to improved reliability, consistency, and confidence in radiologic practices. In addition, it provides new perspectives on the development of explanation methods not solely relying on visualization maps Papanastasiou et al. (2023). A breast cancer classification task was proposed as a case study to evaluate the effectiveness of Rad4XCNN. To this end, we acquired two in-house datasets, one for training and internal testing and the other for external validation.
@@
  • the use of deep architectures for breast cancer classification in ultrasound imaging exploiting an online dataset, and two in-house datasets for internal test and external validation;
@@ 2 Background
The proposed method embraces the explainability problem by exploiting the advantages and disadvantages of deep learning and radiomic workflow-based classification methods. Therefore, the section introduces general concepts related to XAI and particularizes deep architectures and radiomics. At the end of the section, the need and importance of the proposed method are justified.
2.1 Explainable AI
@@
Numerous definitions of explainable AI are not fully standardized yet Longo et al. (2024). For this article, it is essential to delineate certain concepts. A model is intrinsically interpretable when its transparent structure allows an understanding of its decision process. For this reason, a transparent algorithm doesn’t require explanation methods for its interpretation. From an algorithmic point of view, Logistic Regression, Decision Tree, and Naive Bayes are considered transparent, while Neural Networks-based methods are defined as black-boxes. In the case of black-box algorithms, explanation methods are required for their introspection. An explanation can be global and local. A global explanation focuses on explaining a model’s behavior across its entire dataset. A local explanation explains the prediction of a specific dataset sample (i.e. patient). To an intelligible feature it is possible to associate a human-understandable meaning. Radiomic features are considered intelligible. On the other hand, deep features that are extracted using neural network-based architectures, lack interpretability (are not intelligible). We can call features extracted via convolutional neural networks CNN-derived, and extracted via Vistion Transformer as ViT-derived.
@@
The combination of intelligible features and transparent algorithms makes the whole system intrinsically interpretable for two reasons: 1) it is possible to calculate the impact of the feature for model decision; 2) it is possible to correlate and validate the model’s findings comparing it with the clinical literature. Local and global explanations are possible by employing black-box algorithms and intelligible features. These techniques are usually implemented to the already trained model and are known as post-hoc algorithms. When not-intelligible features are used, model explainability becomes significantly more complex: model findings cannot be compared with clinical literature because each feature’s meaning is unknown. A method is defined agnostic when can be applied to any algorithm and architecture. For classification tasks, an XAI method can be class-dependent when explanations are given for each class separately, conversely, they are defined as class-independent.
2.2 Explainability in Deep Architectures
@@
Convolutional-based neural networks have emerged as a standard for deep feature extraction in medical images. Deep learning has shown outstanding capabilities in diagnostic imaging across various diseases and modalities, highlighting its potential as a valuable clinical tool. Despite this promise, its adoption in clinical settings remains limited. One of the main reasons is related to the lack of transparency and trust de Vries et al. (2023). The idea of saliency maps, which is used to highlight the visual regions that are most important for the prediction, is the primary emphasis of these architectures’ explainability Itti et al. (1998); Mamalakis et al. (2023). In Simonyan et al. (2013), given an input image, the gradient computation of the class score is calculated to visualize the activation map for a particular class. Integrated gradients are proposed in Sundararajan et al. (2017) for a more robust explanation through activation maps, adding the constraint of respecting the ’sensitivity’ and the ’implementation invariance’ principles/axioms. In Zeiler et al. (2011) and Zeiler and Fergus (2014) a network using deconvolution is proposed to visualize the convolutional network while in Springenberg et al. (2014) the guided back-propagation is introduced to visualize the features learned from CNN. The most popular algorithms exploiting saliency belong to the category of ’class activation maps’ (CAM) Zhou et al. (2016a). GradCAM Selvaraju et al. (2017) has been introduced after CAM, to overcome the limitation of being able to visualize only the last layer and only particular CNN architectures. Despite many other methods being proposed later, GradCAM is still one of the most popular. Lately, gradient-free methods are also gaining wide popularity to overcome some limitations for computing gradients in very-deep architectures, and overcome the problem of anomalous behavior in incorrect prediction scenarios Muhammad and Yeasin (2020).
@@
Saliency maps are mainly used to provide local explanations and it is not clear how these tools can be used for global explanation. Furthermore, it is important to emphasize that criticisms have been raised about their use Prinzi et al. (2023a). According to Oh et al. (2020); Signoroni et al. (2021), for instance, poorly localized and spatially blurred vision was discovered in certain instances. Furthermore, it has been demonstrated that various methods for computing saliency maps can yield contradictory outcomes Prinzi et al. (2024a); Zhang et al. (2022).
2.3 Explainability in Radiomics
@@
Radiomics is gaining prominence for extracting features from radiological images. It operates on the premise that some pixel-wise patterns are hidden from the human eye Gillies et al. (2016). In contrast to deep features, radiomic features quantify certain properties like texture/pattern and statistical measurement by applying mathematical formulas to the images. For this reason, radiomic features are also called hand-crafted, counterposed to the learned extracted by neural networks. Radiomic flow has several advantages over deep extraction. Radiomic feature extraction does not require a large amount of data, as is the case with deep feature extraction Wei (2021). Moreover, shallow learning algorithms prove to be well-suited for classifying radiomic data, widely recognized as more appropriate for training with small datasets. The primary benefit of radiomic features is their inherent interpretability, as their meanings are well understood Militello et al. (2023); Prinzi et al. (2024b). Using shallow learning algorithms alongside these inherently interpretable features allows for both local and global explanations of the predictive models.
@@
2.4 Our Rationale
Despite in some applications the use of transparent models can be sufficient Rudin (2019), the design of a transparent model can introduce a level of complexity Chen et al. (2022): deep features tend to be more informative than interpretable radiomic features. In addition, the use of transparent algorithms (such as linear models) may fail when the relationships between the data are complex and nonlinear. As a consequence, the use of transparent-by-design models can lead to poor performance. For this reason, the eternal conflict between the explainability and accuracy trade-off comes into play.
The literature indicates the need for new methods for CNN explanation that deviate from those based on visualization maps Papanastasiou et al. (2023). Some attempts have been made by Paul et al. (2019); Wang et al. (2019); Chen et al. (2016) although there is no explicit reference to XAI. Paul et al. (2019) evaluated how radiomic model performance varies by replacing a correlated deep feature in the radiomic signature. However, regarding the performed interpretable feature extraction, the Image Biomarkers Standardization Initiative (IBSI) guidelines Zwanenburg et al. (2020) had not yet been proposed at that time. As a consequence, assessing the reproducibility of the study and the resulting impact of the study can be challenging. In addition, the authors do not use state-of-the-art deep extraction architectures such as ResNet, DenseNet, and ViT, whose depth is quite different from the VGG employed. Furthermore, they agree with an important limitation regarding the fact that a single slice was used to extract deep features in a volumetric image, whereas semantic information was generated from multiple slices. Above all, in contrast to our idea, they modify the input signature by replacing semantic features with deep ones, introducing a kind of constraint for model training. Wang et al. (2019) attempts to exploit the radiologist’s knowledge to qualitatively and visually correlate CNN-based feature maps with semantic features related to the investigated domain. However, these radiological features are operator-dependent, as is their correlation with the feature maps. As a result, their proof-of-concept is not generalizable to any context or domain. A similar idea was proposed in Chen et al. (2016), to reduce the gap between deep features and various clinical semantic features. The model aims to yield semantic rating scores to support a deeper analysis for either clinical diagnosis or educational purposes. More recently, Liu et al. (2023) tried to train deep models to grade informative semantic features related to the analyzed task rather than directly predicting the disease malignancy. Both papers lack formalism regarding the terminology related to explainable AI (i.e., the characteristics of the method in terms of global-local explanation, applicability, class dependencies, etc.) In general, in the attempt to make the predictive process transparent, the cited articles, employ semantic features subject to inter-operator variability or modify the training process, which could harm the established predictive performance of deep models.
In this paper, we present Rad4XCNN, a method that integrates deep features’ predictive strengths with the interpretability of radiomic features. The method does not introduce any constraints in the training that can inhibit its accuracy. Considering the development of new explanation methods not solely based on visualization maps as a major challenge Papanastasiou et al. (2023), Rad4XCNN offers an innovative approach to explain deep architectures and provides a global explanation of deep networks, an aspect often overlooked in current research.